FEL Upgrade Project Weekly Brief
January 8-12, 2001


Our lattice designer (Dave Douglas) released Version 1.1 of the accelerator driver design to engineering this week.
Congratulations to Dave for achieving this milestone.


Final preparations for next week's Laser Processing Consortium/FEL User's Meeting workshop which will be held at Jefferson
Lab on Jan. 18-19th.

We continued preparation of the cost and schedule documentation for the Phase 2 of the FEL upgrade project which we are
planned to integrate with Phase 1 after the Phase 2 funding has arrived (hopefully next month).

WBS 3 (Beam Physics):

A POISSON model for the gun was resurrected and is available for gun design modification studies.

The pre-release Revision 1.1 machine design was released to WBS 9 for layout and engineering. This version nominally
incorporates the near-concentric optical cavity, resolves various interferences, allows additional space for compaction
management elements, supports injection/extraction line geometry providing the desired beam quality, and uses parts that can be
common with the UV system. It is optimized for the probable initial operating energy of 145 MeV.

WBS 4 (Injector):

The use of the new reverse flow heater on the current wafer gave a uniform QE in the area that has consistently had good QE
but lacked uniformity.  The portion of the wafer with a low QE  failed to recover to a higher level indicating more than a heater

The preliminary design for a cathode shield and internal cesiation scheme has been completed for the upgrade gun design.
Conceptual design for the gun chamber has been presented and awaits review.  The rate monitor was received and installed in
the system for mapping the focused cesiator.  The system is being set up for bake out.

WBS 5 (SRF):

Waveguide tooling design has been completed by the vendor.
Vendor proposal for modification to design has been reviewed and approved.
Thermal shield procurement package has been completed and the purchase request signed.

WBS 6 (RF):

Continuing discussions with Glassman High Voltage on the upgrade of the Gun's high voltage power supply. No progress in the
installation and test efforts for zones 3 and 4 because we did not get a repaired LCW system until the end of the week.

We are pleased to report that he LCW was brought back on line Friday apparently without contamination.

The debug of the RF operational systems was completed for the upcoming run.

WBS 8 (Instrumentation):

The updates to the Laser Safety System (LSS) was completed this week, and all 4 of the operational user labs
were certified. All of the certification procedures have also been updated to reflect the changes to the system -
there is now a key in each lab that enables it to be operated in a "Hutch" mode allowing users to be present
during experiments providing the FEL beam is enclosed and interlocked. This system will also allow for
switchyards to be installed so that multiple hutches can be installed in a given lab with no additional changes to
the LSS. Many thanks to the four operators and to Ron Lauze' for participating in the certifications. This activity
doubled as an eight hour training period so they are now qualified to operate the FEL.

The alignment mode changes are complete and documented. This change will allow the "short pulse" users
macro pulses up to 60 Hz, the previous limit was 2 Hz. The expanded Beam Mode will be certified on the 1/12/01
run. Entry status lamps were added to the control room and to each of the user lab doors. With the expanded
modes of operation these (12) stoplights will make it easy for the FEL users to know when one can go into a lab
and whether goggles are required. Additional video monitors were added in the control room for monitoring the
user labs and mirror cans/ beam dumps.

Work continues on the optical BPMs. There are four prototypes built for this upcoming run, these will be used with
picomotors to create a lock on the optical transport system. The error signals will be used to drive each of the
mirror cans to a zero offset.

Parts continue to be purchased for the upgrade to the vacuum system. Good progress continues with the 3" BPMs and the 3"
shielded beam viewers.

WBS 9 (Transport):

o George Biallas discussed designing the Injection and Extraction dipoles (GU & GV) with DULY Research at their
o Their magnetic analysis showed that the mu metal pole is a magnetic short if carried to the edge of the return leg and
   creates a smaller good field region.  When the mu metal is placed only inside  the coils, the good field region expands
   significantly.  We will search for an optimum point somewhere within the coil bundle where the good field
   region is maximized.
o Their analysis also shows that there is a natural field gradient in the field at the plane perpendicular to the beam at the
   center plane of the magnet.  This means that the magnet is so narrow and the gap so big that the end field drop-offs
   that begin within the magnet gap are intersecting at the narrow end of the magnet.  We discussed closing the
   gap enough at the narrow end of the magnet to cancel the end effect.
o They suggested the use of permanent magnets for this application which a simplistic analysis showed has excellent
   field uniformity.  This could be the basis of an additional task.
o George Biallas discussed designing the First Reverse Bend, (GX), the Second Reverse Bend (GQ) and the 180
   degree bend (GY) with Advanced Energy Systems at their office.
o The most significant issue with the design of the 180 degree dipole is the insertion of the path length correctors into the
   ends.  As specified, at 35,000 G-cm, it is superposing a 1.75 KG dipole (for 20 cm length) on to the ends of a
   magnet that is 7.5 KG already.  This implies adding to the local return leg and pole tip iron thickness by almost 25 %.
   Since this specified field is for the worst case of a 180° phase advance at 210 MeV.  We agreed to clarify how the
   correctors would be used.
o Dave Douglas has formally  issued the "cut iron" engineering version of his beam transport layout to be drawn to
   check for interference.


3 inch bore magnets
We finally received an extensively researched quote on the quadrupole from our principle coils supplier.   While the yoke
quotation was expensive, the coil and assembly quotation were within our estimates.  So far, the use of the best quotations falls
within out estimates.

Trim Quad for the Arcs
Magnetic design work has started.

o Measurement Probe: The fiberglass plastic parts for the probe body are being annealed.

Sextupoles and Octupoles
Magnetic design work has started.

o We ordered the shielded bellows for the vacuum piping.

WBS 10 (Wiggler):

Progress is being made on finalizing the wiggler assembly.

WBS 11 (Optics):

With receipt and initial checkout of the deformable high reflecting mirror design that came out of our contract at
AES, we completed the design of the internal components of the optical cavity mirror test stand. Except for a few
minor corrections, the prints are signed off and will be going out for bid.

We have started the design of the optical cavity assemblies.  So far, it is clear we can switch between three
wavelength ranges.  At this time we are trying to make it an UHV enclosure, so we can carry over the design to
the UV FEL.

We installed other O-BPM optics on the upstream optical table, to monitor the FEL position near the outcoupler,
and near mirror can 2, which is in the vault.  We upgraded the optics at the end-of-line, specifically a larger pickoff
wedge was installed, and we changed the optics for the ModeMaster and O-BPM to more appropriate sizes.
We received 6 of New Focus' OEM single-channel picomotor controllers, the very first released to the public.
They will be integrated into a new chassis for control of the mirrors associated with the O-BPM system.
The vacuum installation group swapped the LSS shutter with the "hole harp", an intracavity aperture for mode
control.  They did a good job, as did the I&C group with the new cabling.  This should be ready for the run.

Work done specifically to prepare for operations included:

Swapping out the 1 micron mirrors for 3 micron mirrors.
Replacing the black-anodized mirror holders in the collimator linear stage with unanodized mirrors.


We prepared for "hot check-out" next week by subsystem check outs this week as noted in the discussions above with the
injector, rf, instrumentation and optical systems.

Privacy and Security Notice